OpenStudy (anonymous):

what the hell Uhh this problems -_- Find the missing term (x^12)^5 * (x^-2)^9 * ? = (x^40)^5

OpenStudy (anonymous):

here is a pic so u can see more clear

OpenStudy (anonymous):

So (x^6)^14 * ? = (x^40)^5

OpenStudy (anonymous):

so ur saying the missing term is (x^6)^14 ..?

OpenStudy (anonymous):

No I just made the equation smaller. Instead of solving (x^12)^5 * (x^-2)^9 * ? = (x^40)^5, you can solve (x^6)^14 * ? = (x^40)^5 equation.

OpenStudy (anonymous):

@timo86m

Nnesha (nnesha):

first of all coefficient should be 1 because answer have one coefficient

OpenStudy (anonymous):

that did nothing

Nnesha (nnesha):

and you have to multiply exponent

OpenStudy (anonymous):

so do i multiply the ones that are inside or outside the parentesis..?

Nnesha (nnesha):

yes

OpenStudy (anonymous):

but how do i get the missing term?

Nnesha (nnesha):

wait still working on it

OpenStudy (anonymous):

alright im here :)

Nnesha (nnesha):

do you have answer choice s

OpenStudy (anonymous):

no i wish :/ ..u can see i post a picture on top..of it is more better if u like to see it

OpenStudy (anonymous):

but foreal this problemis really complicated -_-

Nnesha (nnesha):

well i think it can be x^79)square

OpenStudy (anonymous):

maybe right

Nnesha (nnesha):

because ..... so left side should be equal to 200 exponent